# Covers of surfaces

### v1

@@ 3.2 Non-locally finite covers of infinite-type surfaces
We observe that if $\alpha$ is a separating curve in $S$, such that both sides of $A$ and $B$ of $S\smallsetminus\alpha$ are not precompact, then $\alpha$ can be extended to a free basis for $\pi_{1}(A)$ and $\pi_{1}(B)$. By Van Kampen’s theorem, $\pi_{1}(S)=\pi_{1}(A)*\pi_{1}(B)$. It is not hard to see from this that any basis for $\pi_{1}(Y)$ can be extended to a basis for $\pi_{1}(S)$. ∎
@@
3.3 The universal abelian cover
@@
Theorem 1.2 (Universal abelian covers).
@@
3.4 $\mathbb{Z}/n\mathbb{Z}$-homology covers
@@
Deferring the proof of Lemma 3.9, we now prove prove Theorem 1.3.
@@
The first step is to prove:
Claim 3.10.
$\tilde{S}$ has infinite genus.
Proof.
Since $\pi$ is a regular cover of infinite degree, it suffices to prove that $\tilde{S}$ has positive genus.
@@
If $S$ has genus zero, take some essential $4$-punctured sphere $T\subset S$. Then $S\smallsetminus T$ has four components. Moreover, all four components of $S\smallsetminus T$ are noncompact: any compact component is a planar surface with one boundary component, and hence is a disk, contradicting that $T$ is essential. It follows that $H_{1}(T,\mathbb{Z})\hookrightarrow H_{1}(S,\mathbb{Z})$. Hence, if $\tilde{T}\subset\pi^{-1}(T)$ is a component, then $\tilde{T}\to T$ is the $\mathbb{Z}/n\mathbb{Z}$-homology covering of $T$, which has degree $n^{3}$, so $\chi(\tilde{T})=-2n^{3}$. Each component of $\partial T$ is primitive in $H_{1}(S,\mathbb{Z})$, so unwraps $n$ times in $\tilde{T}$, and hence there are $n^{3}/n=n^{2}$ lifts of each, for a total of $4n^{2}$ boundary components. A quick computation shows that $\tilde{T}$ has genus $g=1-\frac{1}{2}(-2n^{3}+4n^{2})$, which is never zero for $n\geq 2$. ∎
By Lemma 3.9 and Claim 3.10, $\tilde{S}$ is either the Loch Ness monster surface or its spotted version. If $S$ has isolated planar ends, these ends lift to isolated planar ends in $\tilde{S}$, so $\tilde{S}$ is the spotted Loch Ness monster surface. Conversely, since the cover is regular, any isolated planar end in $\tilde{S}$ covers an isolated planar end in $S$. So if $S$ has no isolated planar ends, $\tilde{S}$ is the Loch Ness monster surface. ∎
3.4.1 The proof of Lemma 3.9
Before formally proving the Lemma 3.9, we give a brief heuristic of the proof.
@@
Figure 4: The compact subsurface $K$, the reference bi-infinite arcs $\alpha_{-}$ and $\alpha_{+}$, and the arcs $b_{-}$, $b_{+}$, and $a$ used to form $\gamma$.
@@ 4 Arbitrary covers
Theorem 1.1 (Everything covers everything).
@@
The idea for the proof of Theorem 1.1 is to show that any surface $S$ with non-abelian fundamental group is covered by the blooming Cantor tree surface $\mathcal{B}$ and that any infinite-type surface admits a $\pi_{1}$-injective embedding into $\mathcal{B}$, see Proposition 4.2. Then, Lemma 2.4 gives us Theorem 1.1.
Before starting the proof of Theorem 1.1 we show that the blooming Cantor tree surface regularly covers the Loch Ness monster surface.
@@
Figure 5: The blooming Cantor tree surface with multicurve $\Lambda$ depicted in red.
@@
Figure 6: The once-punctured Loch Ness monster surface as a subsurface of $\mathcal{B}$ according to our construction. The left side of the figure represents the sole planar end, while the right side represents the end accumulated by genus.
Proof of Theorem 1.1.
By Proposition 4.2 and Lemma 2.4 it suffices to show that given $S$ there exists a covering space $B$ homeomorphic to a blooming Cantor tree surface. Since $S$ has non-abelian $\pi_{1}$ we can find a $\pi_{1}$-injective pair of pants $P\subset S$. Then, by Lemma 2.4 we get a cover $S^{\prime\prime}\to S$ where $S^{\prime\prime}$ is an open pair of pants. By Theorem 1.2 $S^{\prime\prime}$ is covered by a Loch Ness monster surface $L$ and we conclude by Lemma 4.1. ∎
Appendix A Appendix
@@ A.1 Group cohomology
$\mathbb{Z}^{G}:=\{\phi\mskip 0.5mu\colon\thinspace G\to\mathbb{Z}\},$
which we consider as an abelian group under addition of functions. We realize $\mathbb{Z}^{G}$ as a $G$-module by defining the (left) $G$-action $(g\phi)(h)=\phi(gh)$. So, the action of $g\in G$ precomposes $\phi$ with left translation by $g$. From this definition we observe the following equality, which will be used several times: if $x\in\mathbb{Z}^{G}$ and $g,h\in G$, then
$(gx)(h)=x(g^{-1}h)$ (1)
@@
where $c_{g}\in\mathbb{Z}$ and—abusing notation—we view $g$ as the indicator function that assigns $1$ to $g$ and 00 to all other group elements. The elements of $\mathbb{Z}G$ can then be realized as the formal sums in which $c_{g}=0$ for all but finitely many g. We will switch back and forth between formal $\mathbb{Z}$-linear combinations and functions as convenient.
@@
Now, 00-chains are functions $\phi:\{1\}\to M$, or in other words, a choice of an element of $M$. By (2), $\partial^{0}\phi=0$ exactly when the element $\phi(1)\in M$ is invariant under the action of $G$, and hence we get
@@
Here are some examples for different $G$-modules $M$.
  1. 1.
@@
To understand $1$-cocycles, first note that if $\phi$ is a $1$-cochain, then
@@
Setting the right-hand side to zero, we see that $\phi$ is a $1$-cocycle if and only if
@@
    where the third equality uses (1) and the second-to-last equality uses the cocycle condition. Therefore, $\phi=\phi_{x}$, implying that $\phi$ is a coboundary. It follows that $H^{1}(G,\mathbb{Z}^{G})=0$.
A.2 Ends via group cohomology
In this section we describe how the set of ends of $G$ is encoded in the cohomology group $H^{1}(G,\mathbb{Z}G)$. It is well known that the number of ends of an infinite finitely generated group $G$ is equal to $1+\mathrm{rank}(H^{1}(G,\mathbb{Z}G))$, where rank denotes the cardinality of a maximal $\mathbb{Z}$-linearly independent subset (see for example [7, Theorem 13.5.5]). However, we require a more transparent proof of this fact that will allow us to more easily generalize to the relative setting discussed in Theorem 2.6.
@@ A.2 Ends via group cohomology
If $g\in\partial x$, then there exists $s\in S$ such that $x(sg)\neq x(g)$. By (1), it follows that $s^{-1}x(g)\neq x(g)$ and hence $s^{-1}x(g)-x(g)\neq 0$, which says that $g$ is an element in the support of $(s^{-1}-1)x$. It follows that $\partial x$ is contained in the union of the supports of the elements $(s-1)x\in\mathbb{Z}^{G}$ for $s\in S$. If we assume that $\phi_{x}$ is a 1-cochain with $\mathbb{Z}G$ coefficients, then $(s-1)x=\phi_{x}(s)\in\mathbb{Z}G$ for all $s\in S$; it follows that $\partial x$ is finite, establishing the forwards direction.
Now, assume $\partial x$ is finite. Fix $g\in G$. If $g^{-1}=s_{1}\cdots s_{m}$, where $s_{i}\in S$, we claim that $(g-1)x$ is supported within the $m$-neighborhood of $\partial x\subset G$, with respect to the word metric. Indeed, if $h$ lies outside this neighborhood, then for each $i$, the elements $s_{i}\cdots s_{m}h$ does not lie in $\partial x$, so $x(s_{i}\cdots s_{m}h)=x(s_{i-1}\cdots s_{m}h)$, and inductively we get $x(g^{-1}h)=x(h)$. Using (1), we have that $gx(h)-x(h)=0$, so $(g-1)x(h)=0$ implying that $\phi_{x}(g)=(g-1)x\in\mathbb{Z}G$. As $g\in G$ was arbitrary, we conclude that $\phi_{x}$ is a 1-cochain with $\mathbb{Z}G$ coefficients. ∎
Lemma A.2 ($1$-cocycles with $\mathbb{Z}G$-coefficients).
@@
By definition, $\phi_{x}$ is a 1-coboundary if and only if there exists $y\in\mathbb{Z}G$ such that $\phi_{x}=\phi_{y}$. For the forwards direction, suppose that $\phi_{x}$ is a 1-coboundary. Then there exists $y\in\mathbb{Z}G$ such that $(g-1)x=(g-1)y$ for every $g\in G$. In particular, $g(x-y)=x-y$ for each $g\in G$, implying that $x-y$ is constant. Hence, $x$ is constant off the finite support of $y$. Conversely, suppose that $x$ is constant outside a finite subset $Y$ of $G$. Let $d\in\mathbb{Z}$ such that $x(g)=d$ for all $g\notin Y$. Define $y\in\mathbb{Z}G$ by $y(g)=0$ if $g\notin Y$ and $y(g)=x(g)-d$ if $g\in Y$. Then $x-y$ is constant, and by repeating the computation above in reverse, $\phi_{x}=\phi_{y}$, implying that $\phi_{x}$ is a 1-coboundary ∎
@@
Next, we show the map is injective: suppose $x\in\mathbb{Z}^{G}$ with finite boundary such that $\bar{x}\in Z$. As $\bar{x}\in Z$, there exists $d\in\mathbb{Z}$ such that $\bar{x}(\xi)=d$ for all $\xi\in\mathcal{E}(G)$. It follows that $x(g)=d$ for every $g$ that lies in an infinite component of $Cay(G)\smallsetminus\partial x$. As $\partial x$ is finite, Lemma A.3 implies that $[\phi_{x}]=0$. Now, we turn to surjectivity. If $f\in C(\mathcal{E}(G),\mathbb{Z})$, then the continuity of $f$ and the compactness of $\mathcal{E}(G)$ imply that the image of $f$ is finite; let $\{d_{1},\ldots,d_{m}\}\subset\mathbb{Z}$ be the image of $f$. The preimages $E_{i}:=f^{-1}(d_{i})$ are clopen sets that partition $\mathcal{E}(G)$; therefore, there exists a finite subset $F$ of $G$ such that each of the $E_{i}$ are the accumulation points in $\mathcal{E}(G)$ of a component $C_{i}$ of $Cay(G)\smallsetminus F$. Define $x\in\mathbb{Z}^{G}$ such that $x(g)=d_{i}$ for all $g\in C_{i}$ and $x(g)=0$ otherwise. The boundary $\partial x$ is contained in the $1$-neighborhood of $F$, so $\partial x$ is finite, and $f=\bar{x}$. By Lemma A.2, $\phi_{x}$ is a 1-cocycle, and by construction, the image of $[\phi_{x}]$ is $[f]$, establishing the surjectivity of the map.
@@
It not relevant to the work that follows, but we note that as $H^{1}(G,\mathbb{Z}G)$ is a free abelian group (see [7, Theorem 13.5.3]), if $\mathcal{E}(G)$ is infinite, then $H^{1}(G,\mathbb{Z}G)$ is isomorphic to a countably infinite direct sum of copies of $\mathbb{Z}$.
@@ A.3 Swarup’s Theorem
By Lemma A.2, any $1$-cocycle with $\mathbb{Z}G$-coefficients is of the form $\phi_{x}$, where $x\in\mathbb{Z}^{G}$ has finite boundary. The map $r$ takes $[\phi_{x}]$ to the class in $H^{1}(H,\mathbb{Z}G)$ of the restriction $\phi_{x}|_{H}:H\to\mathbb{Z}G$ given by

### v3

@@ 3.2. Non-locally finite covers of infinite-type surfaces
Let $\alpha$ be a separating curve in $S$ with complementary components $\Sigma_{1}$ and $\Sigma_{2}$. We observe that if neither $\Sigma_{1}$ nor $\Sigma_{2}$ is precompact, then $\alpha$ can be extended to a free basis for $\pi_{1}(\Sigma_{1})$ and $\pi_{1}(\Sigma_{2})$. By Van Kampen’s theorem,
$\pi_{1}(S)=\pi_{1}(\Sigma_{1})*\pi_{1}(\Sigma_{2}).$
It is not hard to see from this that any free basis for $\pi_{1}(Y)$ can be extended to a free basis for $\pi_{1}(S)$. ∎
@@
3.3. The universal abelian cover
@@
Proposition 1.2 (Universal abelian covers).
@@
3.4. $\mathbb{Z}/n\mathbb{Z}$-homology covers
@@
Deferring the proof of Lemma 3.9, we now prove Theorem 1.3.
@@
We first show that $\tilde{S}$ has infinite genus. Since $\pi$ is a regular cover of infinite degree, it suffices to prove that $\tilde{S}$ has positive genus.
@@
If $S$ has genus zero, take some essential 444-punctured sphere $T\subset S$. Then $S\smallsetminus T$ has four components. Moreover, all four components of $S\smallsetminus T$ are noncompact: any compact component is a planar surface with one boundary component, and hence is a disk, contradicting that $T$ is essential. It follows that $H_{1}(T,\mathbb{Z})\hookrightarrow H_{1}(S,\mathbb{Z})$. Hence, if $\tilde{T}\subset\pi^{-1}(T)$ is a component, then $\tilde{T}\to T$ is the $\mathbb{Z}/n\mathbb{Z}$-homology covering of $T$, which has degree $n^{3}$, so $\chi(\tilde{T})=-2n^{3}$. Each component of $\partial T$ is primitive in $H_{1}(S,\mathbb{Z})$, so unwraps $n$ times in $\tilde{T}$, and hence there are $n^{3}/n=n^{2}$ lifts of each, for a total of $4n^{2}$ boundary components. A quick computation shows that $\tilde{T}$ has genus $g=1-\frac{1}{2}(-2n^{3}+4n^{2})$, which is never zero for $n\geq 2$. Thus, by considering the images of $\tilde{T}$ under the infinite deck group, $\tilde{S}$ must have infinite genus.
We now know that $\tilde{S}$ has infinite genus. By Lemma 3.9, $\tilde{S}$ is either the Loch Ness monster surface or its spotted version. If $S$ has isolated planar ends, these ends lift to isolated planar ends in $\tilde{S}$, so $\tilde{S}$ is the spotted Loch Ness monster surface. Conversely, since the cover is regular, any isolated planar end in $\tilde{S}$ covers an isolated planar end in $S$. So if $S$ has no isolated planar ends, $\tilde{S}$ is the Loch Ness monster surface. ∎
3.4.1. The proof of Lemma 3.9
Before formally proving Lemma 3.9, we give a brief heuristic of the proof.
@@
Figure 4. The compact subsurface $K$, the reference bi-infinite arcs $\alpha_{-}$ and $\alpha_{+}$, and the arcs $b_{-}$, $b_{+}$, and $a$ used to form $\gamma$.
@@ 4. Arbitrary covers
Proposition 1.1 (Everything covers everything).
@@
The idea for the proof of Proposition 1.1 is to show that any surface $S$ with non-abelian fundamental group is covered by the blooming Cantor tree surface $\mathcal{B}$ and that any infinite-type surface admits a $\pi_{1}$-injective embedding into $\mathcal{B}$, see Proposition 4.2. Then, Lemma 2.4 gives us Proposition 1.1.
Before starting the proof of Proposition 1.1 we show that the blooming Cantor tree surface regularly covers the Loch Ness monster surface.
@@
Figure 5. The blooming Cantor tree surface with multicurve $\Lambda$ depicted in red.
@@
Figure 6. The once-punctured Loch Ness monster surface as a subsurface of $\mathcal{B}$ according to our construction. The left side of the figure represents the sole planar end, while the right side represents the end accumulated by genus.
Proof of Proposition 1.1.
By Proposition 4.2 and Lemma 2.4 it suffices to show that given $S$ there exists a covering space $B$ homeomorphic to a blooming Cantor tree surface. Since $S$ has non-abelian $\pi_{1}$ we can find a $\pi_{1}$-injective pair of pants $P\subset S$. Then, by Lemma 2.4 we get a cover $S^{\prime\prime}\to S$ where $S^{\prime\prime}$ is an open pair of pants. By Proposition 1.2 $S^{\prime\prime}$ is covered by a Loch Ness monster surface $L$ and we conclude by Lemma 4.1. ∎
Appendix A Ends, group cohomology, and Swarup’s theorem
@@ A.1. Group cohomology
$\mathbb{Z}^{G}:=\{\phi\mskip 0.5mu\colon\thinspace G\to\mathbb{Z}\}$
denote the set of functions from $G$ to $\mathbb{Z}$, which we consider as an abelian group under addition. Given $\phi\in\mathbb{Z}^{G}$ and $g\in G$, let $g\phi$ be the function defined by
$g\phi:\,G$$\to\mathbb{Z}$
$h$ $\mapsto\phi(g^{-1}h).$
We realize $\mathbb{Z}^{G}$ as a $G$-module via this (left) $G$-action.
@@
where $c_{g}\in\mathbb{Z}$ and—abusing notation—we view $g$ as the indicator function that assigns 111 to $g$ and 0 to all other group elements. In this form, the left $G$ action on $\mathbb{Z}^{G}$ described above is given by
$h\cdot\left(\sum_{g\in G}c_{g}\cdot g\right)=\sum_{g\in G}(c_{g}\cdot hg)$
for $h\in G$. The elements of $\mathbb{Z}G$ can then be realized as the formal sums in which $c_{g}=0$ for all but finitely many g. We will switch back and forth between formal $\mathbb{Z}$-linear combinations and functions as convenient.
@@
Now, 0-chains are functions $\phi:\{1\}\to M$, or in other words, a choice of an element of $M$. By (1), $\partial^{0}\phi=0$ exactly when the element $\phi(1)\in M$ is invariant under the action of $G$, and hence we get
@@
Here are some examples of different $G$-modules $M$.
  1. (1)
@@
To understand 111-cocycles, first note that if $\phi$ is a 111-cochain, then
@@
Setting the right-hand side to zero, we see that $\phi$ is a 111-cocycle if and only if
@@
    where the third equality uses the definition of the action and the second-to-last equality uses the cocycle condition. Therefore, $\phi=\phi_{x}$, implying that $\phi$ is a coboundary. It follows that $H^{1}(G,\mathbb{Z}^{G})=0$.
A.2. Ends via group cohomology
In this section we describe how the set of ends of $G$ is encoded in the cohomology group $H^{1}(G,\mathbb{Z}G)$. It is well known that the number of ends of an infinite finitely generated group $G$ is equal to $1+\mathrm{rank}(H^{1}(G,\mathbb{Z}G))$, where rank denotes the cardinality of a maximal $\mathbb{Z}$-linearly independent subset (see for example [10, Theorem 13.5.5]). However, we require a more transparent proof of this fact that will allow us to more easily generalize to the relative setting discussed in Theorem 2.6.
@@ A.2. Ends via group cohomology
If $g\in\partial x$, then there exists $s\in S$ such that $x(sg)\neq x(g)$. It follows that $s^{-1}x(g)\neq x(g)$ and hence $s^{-1}x(g)-x(g)\neq 0$, which says that $g$ is an element in the support of $(s^{-1}-1)x$. It follows that $\partial x$ is contained in the union of the supports of the elements $(s-1)x\in\mathbb{Z}^{G}$ for $s\in S$. If we assume that $\phi_{x}$ is a 1-cochain with $\mathbb{Z}G$ coefficients, then $(s-1)x=\phi_{x}(s)\in\mathbb{Z}G$ for all $s\in S$; it follows that $\partial x$ is finite, establishing the forwards direction.
[…]
Lemma A.2 (111-cocycles with $\mathbb{Z}G$-coefficients).
@@
[…]
@@
[…]
@@
It not relevant to the work that follows, but we note that as $H^{1}(G,\mathbb{Z}G)$ is a free abelian group (see [10, Theorem 13.5.3]), if $\mathcal{E}(G)$ is infinite, then $H^{1}(G,\mathbb{Z}G)$ is isomorphic to a countably infinite direct sum of copies of $\mathbb{Z}$.
@@ A.3. Swarup’s Theorem
By Lemma A.2, any 111-cocycle with $\mathbb{Z}G$-coefficients is of the form $\phi_{x}$, where $x\in\mathbb{Z}^{G}$ has finite boundary. The map $r$ takes $[\phi_{x}]$ to the class in $H^{1}(H,\mathbb{Z}G)$ of the restriction $\phi_{x}|_{H}:H\to\mathbb{Z}G$ given by